OpenStudy (anonymous):

I'm not sure if i am allowed to post 10 questions at 1 time, but i did. If im not supposed to im sorry. :) Anyway, i need help with these. Thanks so much! How many terms are in the expansion (a + b)^15. Expand (2x - 3y)^3 Expand (2x - 5y)^5 Expand (n - 4)^6 Expand (x - 3)^5 How many terms are in the expansion (x + y)^10 What is the sixth term of (a - b)^7? What is the sixth term of (x + y)^8 The first terms of (x + y)5 is y5 (True or False) What is the fourth term of (2x + 3y)^6

OpenStudy (kinggeorge):

While there's no rule saying so, it would be better to post each question separately.

OpenStudy (anonymous):

ok, thanks for the tip

OpenStudy (anonymous):

I think the binomial theorem. (a+b)^n = SUM (n choose k)a^(n-k) b^k where k = 0 to n should help for all those.